ROTATIONALLY RESOLVED STRUCTURE IN THE FIFTH AND SIXTH TORSIONAL STATES OF AC-ETALDEHYDE: INTERNAL ROTATION ABOVE THE TORSIONAL BARRIER

YUNG-CHING CHOU, CHENG-LIANG HUANG, \underline{I-CHIA CHEN}, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 300, Republic of China; CHI-KUNG NI, A. H. KUNG , Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei, Taiwan 106, Republic of China.

Fluorescence excitation spectrum of acetaldehyde in its transition Ã^{1}A′′−X̃^{1}A′ is analyzed for the torsional states exceeding the barrier. States with torsional vibrational quantum number v_{t} = 5 A + Eand 6 A at term energy 660−927 cm^{−1} are assigned. This region is about 100−340 cm^{−1} above the top of torsional barrier. These states lie between limits of the torsional vibrational motion and a free rotor; the states 5A and 6A lie close to mix and the K states in E sublevel split further. Avoided crossings for ΔK = 0 and Δm = 0 mod 3 states are observed between K = 2 of 14^{0+}15^{4} and 14^{0+}15^{5} and of 14^{0−}15^{4} and 14^{0−}15^{5}. Consequently the K structure of 14^{0+}15^{5} deviates significantly from the expected parabolic shape. Between the inversion doublets 14^{0+} and 14^{0−} the K structures and the K state at the lowest energy are quite different.